OpenStudy (george.g101):

Write the equation of the line that passes through (3, –2) and has a slope of 4 in point-slope form. A.) y + 2 = 4(x – 3) B.) y – 3 = 4(x + 2) C.) x – 3 = 4(y + 2) D.) x + 2 = 4(y – 3)

OpenStudy (george.g101):

Please help @hartnn @perl

hartnn (hartnn):

the equation of line with slope = m and point = x1,y1 is \(y-y_1 = m (x-x_1)\) here, m = 4, x1, y1 = 3,-2 just plug in!

OpenStudy (queenie):

but pay attention to that -2

OpenStudy (george.g101):

Ohhh ok, but there's only one set of points given?

hartnn (hartnn):

you need only one. x1 = 3, y1 = -2 y - (-2) = 4 (x - 3)

OpenStudy (george.g101):

Oh I see, thank you! :)

hartnn (hartnn):

welcome ^_^